OpenStudy (allymariex3):

Solve for x: (x - 9)2 =1 Completing the square.

hero (hero):

No need to complete the square for this one.

hero (hero):

(x - 9)^2 =1 x - 9 = ±sqrt(1) x = ±sqrt(1) + 9

OpenStudy (anonymous):

already its complete square form hence (x-9)^2=1 or (x-9)^ -1=0 or (x-9-1)(x-9+1)=0 (x-10)(x-8)=0 or x-10=0 or x-8=0 hence x=10,8

OpenStudy (anonymous):

x=5

hero (hero):

(x - 9)^2 =1 x - 9 = ±sqrt(1) x = ±sqrt(1) + 9 x = -1 + 9 = 8 x = 1 + 9 = 10

OpenStudy (anonymous):

x=10, 8 oops

OpenStudy (allymariex3):

@Hero thank you :)